OpenStudy (anonymous):

]If p=2 , what is the value of 4p-8+p with small 3 in corner ? Show your work.

OpenStudy (anonymous):

do you have answers

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4times 2 - 8 + 2^3 solve this

OpenStudy (anonymous):

this is your work

OpenStudy (anonymous):

im not going to just do it for you

OpenStudy (anonymous):

i know one sec

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

or 8?

OpenStudy (anonymous):

which do you think

OpenStudy (anonymous):

how did u get 6

OpenStudy (anonymous):

4*2 = 8 -8 = 0 + 2 = 2 * 2 * 2 = 8

OpenStudy (anonymous):

so 8?

OpenStudy (anonymous):

yes 8 do you understand my work

OpenStudy (anonymous):

yes 4x2=8-8=0=2=2x2=8

OpenStudy (anonymous):

you forgot a 2 = 2*2*2

OpenStudy (anonymous):

wait nvm

OpenStudy (anonymous):

4x2=8-8=0=2x2x2=8

OpenStudy (anonymous):

yes